I need help with this. I am thoroughly confused as to what is being asked:
Evaluate the functions for the values of x given as 1,2,4,8, and 16. Rank in order from fastest to slowest the rate at which each function changes with increasing values of x.
a) f(x)=2x-5
b) f(x)=x^2-4x+3
c) f(x)=x^3+4x^2-2x-3
d) f(x)=7^x
e) f(x)=log x

Answer by Alan3354(69443)

a) f(x)=2x-5
Sub 1, 2, 4, 8 and 16 for x
f(1) = 2*1 - 5 = -3
f(2) = 2*2 - 5 = -1
f(4) = 2*4 - 5 = +3
f(8) = 2*8 - 5 = +11
f(16) = 2*16 - 5 = +27
----------------
Do the same for the others.
b) f(x)=x^2-4x+3
c) f(x)=x^3+4x^2-2x-3
d) f(x)=7^x
e) f(x)=log x
-------------
Graphs are handy for evaluating as asked.
dl the FREE software to graph these at
www.padowan.dk.com/graph/
-----------
email me via the thank you note with any questions or to check your work.